Embedded Systems

Assignment 4

Jan Hendrik Dithmar

2031259

Pascal Gwosdek

2505221

Assignment 4  Dithmar (2031259), Gwosdek (2505221)

We did Exercise 1, of course. For exercise 2, we combined the first three parts in one result and the last part in another result, because there was a change in the way we handeled time.

Exercise 2, parts 1, 2 and 3:

Exercise 2, part 4: